OpenStudy (anonymous):

n a certain county, the number of charter schools is 8 less than twice the number of alternative schools. We know that there are 36 charter schools in the county. How many alternative schools are in the county? 44 22 14 30

OpenStudy (purplemexican):

36+8=44 44/2=22 22*2-8=36

OpenStudy (anonymous):

thats not even an answer

OpenStudy (anonymous):

nvm

OpenStudy (purplemexican):

22 is the answer i was giving you the math behind it